Thursday 11/01/12

9:30 A – 7:30P, 8:30P- 12:00A

  • Software install on STORM2.  Hal can’t detect drivers for national instruments panel.
  • Overview of software install: see post
Fly work
  • Collect virgins
  • Sort and cross Espl/sim flies
  • Clear MTDs into 2 new bottles, move existing bottles to virgin collection
  • esc flies from Vivek stocks arrived today, thankyou Kasia.  Moved to stocks for expansion.
  • Order fly stocks
    • 1411:  ph[503] — affects both coding units at ph locus. similar to deficiency
    • 2406 Scm[D1]
    • 1728 Pc1/TM1
    • 4432  TM3, Sb, Ubx-LacZ (should get an anti-Bgal).
    • 3516 hs-en (cDNA).  is en expression maintained from the native locus?
    • 31611 TRiP Psc (does give phenotype)
  • Fix embryos
  • collect PM virgins

embryo DNA FISH

  • preliminary test of post washed Ubx-cy5 embryos shows no evident staining on turnkey.
  • counterstain with Hoechst and WGA-488
  • labeled cent-dig with rb-K27ac, m-Dm01, shp a-dig
  • # check on confocal

STORM imaging of s2 cells

  • 2D imaging in glox, 40,000 frames H3 still lots of switching (it’s all H3 in the cell all in the same view (well from a optical TIRF section), so this probably isn’t going to stop.
  • switching is extremely dense in 647, first 1000 frames have substantial amount of overlapping emission. Try this on multi-fitter?  Also need a higher threshold — some non-switching background spots (these should help the drift correction though).   About frame 15K hitting single molecule emission more often (maybe 20K).  Still no 405.
  • clean over 100,000 frames, still no activator, all spontaneous switching
  • 200,000 frames, still no activator needed.  suspect switching looks sluggish?
  • 561 extremely sluggish switching, suspect laser power issue, must check.  Okay, speeding up a bit after 10,000 frames.  Probably buffer rundown, been driving glox pretty hard, container is completely open
  • 488 okay switching?  some reasonable background, probably better than embryonic tissue though.  should do a double label control or something to test this better.
  • Powers; 405 = .88mW, 488=100mW, 561=90mW 657=130mW, 750=60mW
  • bead calibration.  new bead slide, 1.7 vis to 1 uL IR beads.  A little too dense on the vis, try 1.5 to 1. 1:20,000 a little dense.  Try 1:10,000 and 1:15,000
  • make pyranose oxidase for longer switching. 2.9 mg in 50uL im buffer + 50uL catalase for 100x stock
  • Move camera back.  Realign exit path optics: using 512-quadview bead field center for brightness and focus.
  • 3 rounds of automated runs using parameters from first run.  Try Pyranose — seems to be switching better, should figure out some way to compare in parallel. (or add to identical samples, seal, and check after longer period of time).
  • Microscope off, 11:45PM




This entry was posted in Summaries and tagged , , . Bookmark the permalink.